# Effective synchronization amid noise-induced chaos

### v1

@@ IV.1 Proof of effective synchronization
We now test the extent to which agent ${\mathrm{A}}$’s $\varphi_{{\mathrm{f}},{\mathrm{A}}}^{(k)}$ and ${\mathrm{B}}$’s $\varphi_{{\mathrm{f}},{\mathrm{B}}}^{(k)}$ agree. We use an extensive simulation of the phase map of Eq. (2) with $\Lambda=0.141$ and $N=500$. We extend the simulation depicted in Fig. 2 to a large number of kicks, $k=10^{4}$. For every $k$, using the procedure of Sec. A.2 (Eq. (7)), we find $\varphi^{(k)}_{{\mathrm{f}},{\mathrm{A}}}$ and $\varphi^{(k)}_{{\mathrm{f}},{\mathrm{B}}}$. To quantify the success of our scheme for identifying a common fiducial phase, we consider the statistics of the deviation $\Delta\varphi_{\mathrm{f}}=\varphi_{{\mathrm{f}},{\mathrm{A}}}-\varphi_{{\mathrm{f}},{\mathrm{B}}}$. There are indeed $4,987$ positive and $5,013$ negative $\Delta\varphi_{\mathrm{f}}$ values, both exhibiting identical histograms (the KLD among the two $\ln(\pm\Delta\varphi_{\mathrm{f}})$ distributions is $\mathcal{O}(10^{-3})$), so we only show $|\Delta\varphi_{\mathrm{f}}|$ henceforth. Likewise, in light of the convergence we observe in Sec. III, we shall only plot $S=(S_{\mathrm{A}}+S_{\mathrm{B}})/2$, where $S_{\mathrm{A}}$ and $S_{\mathrm{B}}$ only differ by numerical entropy estimation error $\sim 0.2$ for large $k\gg K_{\mathrm{m}}$.
@@
[…]
Repeating the above for $\Lambda=0.101,0.141,\ldots,0.501$ reveals that the entropy is still exponentially distributed with a (negative) mean that increases with $\Lambda$, asymptoting to 00 for large $\Lambda$, which is consistent with Ref. [21]. For all $\Lambda$’s tested and for $N=125,250,500$, all the corresponding plots fully overlap with Fig. 5(b) (except for the statistically insignificant tails $\ell<10^{-2.5}$).
To conclude, this section has provided concrete evidence supporting the proposal made in Sec. I. Identical nonlinear oscillators exposed to identical noise can enable independent agents to act in concert and perform time-based communications as though their clocks were synchronized, even though they are not and their phases are only statistically determined.
@@ V Discussion
In this paper, we showed that noise-induced synchronization can be extended beyond its recognized limits. Our study demonstrated effective synchronization under noise-driven dynamics with positive Lyapunov exponents, which precludes conventional noise-induced synchronization. This effective synchronization arose from two key features of our dynamics. First, the statistical distribution of phases at a given moment becomes independent of the initial distribution for every realization of waiting times, as noted in Sec. III. Thus, the distributions seen by two independent agents become equal to each other while varying (together) strongly with more kicks. Second, concentrating on dynamics whose Lyapunov exponent is small in magnitude, we found in Sec. IV that the prevalence of low-entropy distributions $q(\varphi)$ permits accurate effective synchronization.
In this section, we assess the potential impact of these results. First, we note the narrow though significant scope of our explicit study. We argue that this generalized synchronization should occur generically for oscillators subjected to non-synchronizing noise, despite possible degrading effects we ignored. We then discuss the power-law trade-off we observed between the desirable rapid convergence and the obtainable precision of the generalized synchronization. Finally, we consider how generalized synchronization might be relevant for explaining forms of cooperative behavior in organisms.
@@ V.1 Limitations
Our demonstration of effective synchronization was made in a narrow context. Among the various types of noise treated in the literature [15, 22], we considered only impulsive noise whose effect could be described by a phase map [19, 21]. Furthermore, the bulk of our study used a simple class of such maps — a cubic polynomials; other phase maps may not culminate in low-entropy distributions or a meaningful fiducial phase. Additional aspects we did not address include counter-effects that oppose synchronization, such as a disruptive noise that acts independently on each oscillator [17] and inevitable differences between the two agents’ oscillators [16]. These degrading effects would surely reduce the degree of effective synchronization that we demonstrated, and may even render our procedure to determine an effective synchronized phase unstable.
@@
These limitations mean that our example is not immediately applicable to realistic conditions. Still, we may argue that the effect is somewhat general and robust. The effective synchronization persists under continuous variation of the phase map $\psi(\varphi)$ through cubic and quintic phase maps and over a range of Lyapunov exponents $\Lambda$. Further, we recover the effective phase $\varphi_{\mathrm{f}}$ using standard-precision computations and conventional sampling methods. This is not a delicate effect analogous to time-reversing a chaotic trajectory. Moreover, all of our results in Sec. IV show a common variation with the distribution entropy $S$, $\Delta\varphi_{\mathrm{f}}\sim e^{S}/\sqrt{N}$, under a range of underlying dynamics that gave rise to the distribution. Under general noise, one may define the Lyapunov exponent [15, 18], the distribution of $\varphi$ values $q(\varphi)$ at a given time, the notion of independence of this $q(\varphi)$ on initial conditions, and the determination of a fiducial phase $\varphi_{\mathrm{f}}$ from a given $q(\varphi)$. Thus, the synchronization we found in our narrow context should plausibly have a counterpart for clocks subjected to more general forms of noise, such as sensory neural networks [28].
[…]
@@
A clear practical limitation of our procedure was the laborious calculations it required. Partly, this labor was due to our choice to demonstrate precise synchronization, i.e., small $\Delta\varphi_{\mathrm{f}}$. This required using phase maps that produce low-entropy distributions (small positive $\Lambda$). This small $\Lambda>0$ implies slow convergence of $q(\varphi)$, as shown in Sec. III. (For our example of Sec. IV, the convergence number $K_{\mathrm{m}}$ was about $10^{2}$; knowledge of the last $\sim 10^{2}$ was needed in order to infer the converged $\varphi_{\mathrm{f}}$.) If instead we had been content to show synchronization to only a few bits of precision, we could have relaxed our requirement of small entropy and small $\Lambda>0$. Then, the large number of kicks needed to obtain a converged $\varphi_{\mathrm{f}}$ would be correspondingly reduced, thus reducing the required computation. We now consider further this tradeoff between computational labor and precision.
@@ V.2 Convergence rate scaling
The central feature of our dynamics that makes effective synchronization possible is the property of convergence, as discussed in Sec. III; the final distribution $q(\varphi)$ depends only on the recent history. On the other hand, if the history required is long, one must take account of many kicks in order to determine the needed converged distribution — a lengthy process. Thus, it is important to know what governs the speed of convergence.
Our empirical study showed a simple behavior of the convergence though $K_{\mathrm{m}}$ — the number of kicks required for convergence to a common $q(\varphi)$. The qualitative trend was as expected: Larger (positive) $\Lambda$ (thence larger typical entropy $S$) gave faster convergence. Our study showed an apparent powerlaw relation between $K_{\mathrm{m}}$ and $\Lambda$, viz., $K_{\mathrm{m}}\sim\Lambda^{-3/2}$. This robust behavior reinforces the overall consistency of our effective synchronization finding. From this formula, one can readily infer the increase in $\Lambda$ needed to effect a desired reduction in $K_{\mathrm{m}}$. Then, from $S\sim 1/\Lambda$ of Ref. [21] and $\Delta\varphi_{\mathrm{f}}\sim e^{S}/\sqrt{N}$ from Sec. IV, one can infer the expected loss of precision in $\varphi_{\mathrm{f}}$. It will be useful to demonstrate the scaling of $K_{\mathrm{m}}$ more generally and extensively.
@@ V.3 Biological implications
We posit that this synchronization scheme could enable biological agents — cells, organelles, or gene expression networks — to cooperate in a wider range of noisy environments than previously envisioned. For example, our mechanism may be relevant to a long-standing experimental puzzle about information transmission between neurons via sequences of pulses or spikes [42]. The timing of these spikes is believed to carry that information [4]; thus, there needs to be some means of calibrating this timing. The present work suggests that the ambient electrical noise in the brain or variations of the sensed environment [5, 6, 28] could aid this calibration. In particular, these effects may serve as sources for common noise, which can produce a generalized notion of synchronization.
@@ Acknowledgements.
We benefited from discussions with Efi Efrati and Colin Scheibner. We thank Chuanyi Wang for providing the efficient functional form of the quintic maps. We also thank Gil Ariel, Omer Granek, Andrej Košmrlj, Muhittin Mungan, Colin Scheibner, and Tony Song for helpful comments on the manuscript. B.S. acknowledges support from the Princeton Center for Theoretical Science. B.S. would like to thank the members of the James Frank Institute of the University of Chicago for their warm hospitality, where the research was initiated.
@@ A.1 Binless estimation methods using samples
(and the same for $q_{\mathrm{B}}^{(k)}$), where $\{\tilde{\varphi}^{(k)}_{{\mathrm{A}}n}\}$ are the instantaneous phase samples of agent ${\mathrm{A}}$, $\{\varphi^{(k)}_{{\mathrm{A}}n}\}$, arranged in increasing order on a ring ($\tilde{\varphi}^{(k)}_{{\mathrm{A}}n+1}>\tilde{\varphi}^{(k)}_{{\mathrm{A}}n}$, with a mod $\mathrm{mod}\,1$ constraint) and $1/N$ ensures normalization. In Sec. IV, where a statistically significant estimate of $q^{(k)}(\varphi)$ was needed, we used, rather, the kernel-density estimate of Eq. (11).
@@
There are two technical limitations involving the above binless estimation methods. First is our limited sample size $N=500$, leading to an errorbar of ±$\pm 0.2$ for entropy values (identified from samples of uniform distribution) and ±$\pm 0.05$ for KLD values (identified from two sets of samples of the same Gaussian distribution). Second is the limited machine precision [21]; the phases are distinguishable only up to $\Delta_{{\mathrm{A}}n\leftarrow{\mathrm{B}}}^{(k)}>2^{-52}$ using the conventional $64\mathrm{bit}$ representation of real numbers. Each sample with $\Delta_{{\mathrm{A}}n\leftarrow{\mathrm{B}}}^{(k)}<10\times 2^{-52}$ is ignored, and we replace $N\to N-1$.
The loss of samples to machine precision is a prevalent limitation. In the long realization ($k=10^{4}$) of Sec. IV, $407$ and $439$ samples remained in the end for ensembles ${\mathrm{A}}$ and ${\mathrm{B}}$, respectively, which means that our estimates are still statistically meaningful. On the other hand, this implies that the low-entropy distributions (which are those that usually culminate in loss of samples) did not appear much; see Fig. 5, where we see no distribution with $S<-21$. Upon continuing this simulation, there were no longer any distinguishable phase samples past $k=1.2\cdot 10^{4}$, at which point the dynamics have attained entropies between $-34<S<-21$. Were lower-entropy distributions to appear, our results of the main text would improve further, as they are particularly useful for having a small discrepancy $\Delta\varphi_{\mathrm{f}}\sim e^{S}/\sqrt{N}$.

### v2

@@ IV.1 Proof of effective synchronization
We now test the extent to which agent $\mathrm{A}$}’s $\varphi_{{\mathrm{f}},{\mathrm{A}}}^{(k)}$ and $\mathrm{B}$}’s $\varphi_{{\mathrm{f}},{\mathrm{B}}}^{(k)}$ agree. We use an extensive simulation of the phase map of Eq. (2) with $\Lambda=0.141$ and $N=500$. We extend the simulation depicted in Fig. 2 to a large number of kicks, $k=10^{4}$. For every $k$, using the procedure of Appendix A.2 (Eq. (7)), we find $\varphi^{(k)}_{{\mathrm{f}},{\mathrm{A}}}$ and $\varphi^{(k)}_{{\mathrm{f}},{\mathrm{B}}}$. To quantify the success of our scheme for identifying a common fiducial phase, we consider the statistics of the deviation $\Delta\varphi_{\mathrm{f}}=\varphi_{{\mathrm{f}},{\mathrm{A}}}-\varphi_{{\mathrm{f}},{\mathrm{B}}}$. There are indeed $4,987$ positive and $5,013$ negative $\Delta\varphi_{\mathrm{f}}$ values, both exhibiting identical histograms (the KLD among the two $\ln(\pm\Delta\varphi_{\mathrm{f}})$ distributions is $\mathcal{O}(10^{-3})$), so we only show $|\Delta\varphi_{\mathrm{f}}|$ henceforth. Likewise, in light of the convergence we observe in Sec. III, we shall only plot $S=(S_{\mathrm{A}}+S_{\mathrm{B}})/2$, where $S_{\mathrm{A}}$ and $S_{\mathrm{B}}$ only differ by numerical entropy estimation error $\sim 0.2$ for large $k\gg K_{\mathrm{m}}$.
@@
Overall, both panels of Fig. 5 are most encouraging, as a common fiducial phase is established within uncertainty much less than $1$. Indeed the most abundant instances surrounding the black point ($\ell\simeq 1.0$, $S\simeq-2.0$) have a phase discrepancy of order $|\Delta\varphi_{\mathrm{f}}|=1\cdot e^{-2}/\sqrt{500}=6.1\cdot 10^{-3}$, which is a fraction of a percent synchronization accuracy. Further, much smaller deviations are abundant: From Fig. 5(b), $10\%$ of the distributions produce reduced discrepancies smaller than $\ell=10^{-1}$, or $|\Delta\varphi_{\mathrm{f}}|=6.1\cdot 10^{-4}$. The fitted error function ($\mathrm{erf}(x)=(2/\pi)\int_{0}^{x}dte^{-t^{2}}$) suggests that the discrepancy predominantly follows a normal distribution. This supports our proposed scaling $|\Delta\varphi_{\mathrm{f}}|\sim e^{S}/\sqrt{N}$, as indeed the predominant peak typically contributes most to the entropy, so the local width is comparable to $e^{S}$ and contains most of the phase samples ($\mathcal{O}(N)$), so the standard deviation in the estimation of the mean decays as $1/\sqrt{N}$.
Repeating the above for $\Lambda=0.101,0.141,\ldots,0.501$ reveals that the entropy is still exponentially distributed with a (negative) mean that increases with $\Lambda$, asymptoting to 0 for large $\Lambda$, which is consistent with Ref. [24]. For all $\Lambda$’s tested and for $N=125,250,500$, all the corresponding plots fully overlap with Fig. 5(b) (except for the statistically insignificant tails for $\ell<10^{-2.5}$).
To conclude, this section has provided concrete evidence supporting the proposal made in the Introduction. Identical nonlinear oscillators exposed to identical noise can enable independent agents to act in concert and perform time-based communications as though their clocks were synchronized, even though they are not and their phases are only statistically determined.
@@ V Discussion
In this paper, we showed that noise-induced synchronization can be extended beyond its recognized limits. Our study demonstrated effective synchronization under noise-driven dynamics with positive Lyapunov exponents, which precludes conventional noise-induced synchronization. This effective synchronization arose from two key features of our dynamics. First, the statistical distribution of phases at a given moment becomes independent of the initial distribution for arbitrary realization of waiting times, as noted in Sec. III. Thus, the distributions seen by two independent agents become equal to each other while varying (together) strongly with more kicks. Second, concentrating on dynamics whose Lyapunov exponent is small in magnitude, we found in Sec. IV that the prevalence of low-entropy distributions $q(\varphi)$ permits accurate effective synchronization.
In this section, we assess the potential impact of these results. First, we note the narrow though significant scope of our explicit study. We argue that this generalized synchronization should occur generically for oscillators subjected to non-synchronizing noise, despite possible degrading effects we ignored. We then discuss the power-law trade-off we observed between the desirable rapid convergence and the obtainable precision of the generalized synchronization. Finally, we consider how generalized synchronization might be relevant for explaining forms of cooperative behavior in living systems.
@@ V.1 Limitations
Our demonstration of effective synchronization was made in a narrow context. Among the various types of noise treated in the literature [18, 25], we considered only impulsive noise whose effect could be described by a phase map [22, 24]. Furthermore, the bulk of our study used a simple class of such maps — a cubic polynomials; other phase maps may not culminate in low-entropy distributions or have a meaningful fiducial phase. These limitations mean that our example is not immediately applicable to realistic conditions. Still, we may argue that the effect is somewhat general and robust. The effective synchronization persists under continuous variation of the phase map $\psi(\varphi)$ through cubic and quintic phase maps and over a range of Lyapunov exponents $\Lambda$. Further, we recover the effective phase $\varphi_{\mathrm{f}}$ using standard-precision computations and conventional sampling methods. This is not a delicate effect analogous to time-reversing a chaotic trajectory. Moreover, all of our results in Sec. IV show a common variation with the distribution entropy $S$, $\Delta\varphi_{\mathrm{f}}\sim e^{S}/\sqrt{N}$, under a range of underlying dynamics that gave rise to the distribution. Under general noise, one may define the Lyapunov exponent [18, 21], the distribution of $\varphi$ values $q(\varphi)$ at a given time, the notion of independence of this $q(\varphi)$ on initial conditions, and the determination of a fiducial phase $\varphi_{\mathrm{f}}$ from a given $q(\varphi)$. Thus, the synchronization we found in our narrow context should plausibly have a counterpart for clocks subjected to more general forms of noise, such as sensory neural networks [31].
In addition, we did not address counter-effects that oppose synchronization, such as direct interactions [47, 48, 49], the presence of an additional degrading noise that acts independently on each oscillator [20, 19], and inevitable differences between the two agents’ oscillators [19]. These must be included to reliably model, e.g., ecological systems where spatial variations in the environment or inherent differences in the individuals are of interest. We comment, however, that these counter-effects plague standard noise-induced synchronization ($\Lambda<0$) just as they would our effective synchronization ($\Lambda>0$). For $\Lambda<0$, it was shown that these effects ultimately lead to an asymptotic $q^{(k)}(\varphi)$ whose nonzero width is determined by the magnitude of the corresponding counter-effect [19, 20] (as opposed to an ever-decreasing width of $q^{(k)}(\varphi)$ for standard noise-induced synchronization). We expect a similar conclusion for $\Lambda>0$, whereby the scaling $\sim e^{S}/N^{1/2}$ might worsen as a function of degrading noise’s magnitude. This will be explored in a future work.
Although our prescription for identifying $\varphi_{\mathrm{f}}$ is somewhat ungainly, the result is about as precise as one might hope. The statistical fluctuations of $\Delta\varphi_{\mathrm{f}}$ appear consistent with an $N$-sample average of a Gaussian distribution whose variance is of order $N\langle\Delta\varphi_{\mathrm{f}}^{2}\rangle$, where $\langle\cdots\rangle$ denotes the average over samples. The entropy of the $\Delta\varphi_{\mathrm{f}}$ distribution is thus of order $\ln(\langle\Delta\varphi_{\mathrm{f}}^{2}\rangle^{1/2})+\log N$ [46]. Now, the scaling shown in Fig. 5 amounts to saying $\ln(\langle\Delta\varphi_{\mathrm{f}}^{2}\rangle^{1/2})\simeq S-\log N$, i.e., $S$ is comparable to the entropy of $\Delta\varphi_{\mathrm{f}}$. Thus, a much more precise $\varphi_{\mathrm{f}}$ than the observed one would need to contain more information than the source distribution $q(\varphi)$ from which it was derived — a contradiction. Thus, we cannot expect other methods to yield qualitative improvements in precise synchronization compared to $\Delta\varphi_{\mathrm{f}}\sim e^{S}/\sqrt{N}$. [^10] One straightforward example of a fiducial phase includes the first circular mean, whose estimation error follows the central limit theorem, thus scaling as $\sigma/\sqrt{N}$ with $\sigma\sim e^{S}$ (for $\sigma\ll 1$, which are the useful distributions; see, e.g., Ref. [50] on directional statistics). We prefered to use the highest-peak position as the fiducial phase since it is more illustrative of the effect we report compared to the circular mean — that peak’s dominance over all others lies at the core of the effective synchronization. Nevertheless, one can hope for more insightful and efficient calculation methods, which might or might not involve a fiducial phase.
@@
A clear practical limitation of our procedure was the laborious calculations it required. Partly, this labor was due to our choice to demonstrate precise synchronization, i.e., small $\Delta\varphi_{\mathrm{f}}$. This required using phase maps that produce low-entropy distributions (small positive $\Lambda$). This small $\Lambda>0$ implies slow convergence of $q(\varphi)$, as shown in Sec. III. (For our example of Sec. IV, the convergence number $K_{\mathrm{m}}$ was about $10^{2}$; knowledge of the last $\sim 10^{2}$ kicks was needed in order to infer the converged $\varphi_{\mathrm{f}}$.) If instead we had been content to show synchronization to only a few bits of precision, we could have relaxed our requirement of small entropy and small $\Lambda>0$. Then, the large number of kicks needed to obtain a converged $\varphi_{\mathrm{f}}$ would be correspondingly reduced, thus reducing the required computation. We now consider further this tradeoff between computational labor and precision.
@@ V.2 Convergence rate scaling
The central feature of our dynamics that makes effective synchronization possible is the property of convergence; the final distribution $q(\varphi)$ depends only on the recent history of waiting times. Interestingly, we found that the amount of history required had a simple dependence on $\Lambda$; the effective number of relevant kicks, $K_{\mathrm{m}}$, followed a simple power law. This observed power-law dependence merits further discussion.
In general, when an iterated map becomes independent of its initial state, the characteristic kick number for convergence $K_{\mathrm{m}}$ has little relation to $\Lambda$. $\Lambda$ characterizes the separation rate of adjacent points while $K_{\mathrm{m}}$ depends on the merger of separated regions. Nonetheless, we empirically find a systematic variation of $K_{\mathrm{m}}$ with $\Lambda$ and, even more so, that $K_{\mathrm{m}}$ diverges as $\Lambda\to 0$. Characterizing this divergence further showed an apparent near-$3/2$ power-law dependence ranging across two orders of magnitude in $K_{\mathrm{m}}$. This suggests that there is some unexplained interdependence between $K_{\mathrm{m}}$ and $\Lambda$, as we sketched qualitatively in Sec. III. However, we have no evidence for such interdependence beyond the cubic map. Understanding this divergence merits a future study.
We note that the convergence found here is quite unusual. In the paradigmatic dynamics exhibiting mixing [37], any initial condition asymptotically approaches an invariant distribution. Here, since the phase map keeps being randomly applied, there is no invariant distribution $q^{(k\to\infty)}(\varphi)$, and instead every consecutive $q^{(k)}(\varphi)$ differs stochastically from the previous. It is the $q^{(k)}$’s with a different $q^{(0)}$ that become progressively more similar with $k$ for a given $\{\beta^{(k)}\}$. Such a convergence phenomenon was shown to occur in single-particle chaotic trajectories under common noise as well [51], and of course occurs in conventional noise-induced synchronization.
For practical purposes, with a known scaling between $K_{\mathrm{m}}$ and $\Lambda$, one can readily predict the increase in $\Lambda$ needed to effect a desired reduction in $K_{\mathrm{m}}$. Then, from $S\sim 1/\Lambda$ of Ref. [24] and $\Delta\varphi_{\mathrm{f}}\sim e^{S}/\sqrt{N}$ from Sec. IV, one can infer the expected loss of precision in $\varphi_{\mathrm{f}}$.
@@ V.3 Biological implications
We posit that this synchronization scheme could enable biological agents — cells, organelles, or gene expression networks — to cooperate in a wider range of noisy environments than previously envisioned. For example, our mechanism may be relevant to a long-standing experimental puzzle about information transmission between neurons via sequences of pulses or spikes [49]. The timing of these spikes is believed to carry that information [4, 8, 9]; thus, there needs to be some means of calibrating this timing. The present work suggests that the ambient electrical noise in the brain or variations of the sensed environment [5, 6, 31] could aid this calibration. In particular, these effects may serve as sources for common noise, which can produce a generalized notion of synchronization.
@@ Acknowledgements.
We benefited from discussions with Efi Efrati and Colin Scheibner. We thank Chuanyi Wang for providing the efficient functional form of the quintic maps. We also thank Gil Ariel, Eli Barkai, Omer Granek, Andrej Košmrlj, Adilson Motter, Muhittin Mungan, Colin Scheibner, and Tony Song for helpful comments on the manuscript. B.S. acknowledges support from the Princeton Center for Theoretical Science. B.S. would like to thank the members of the James Frank Institute of the University of Chicago for their warm hospitality, where the research was initiated.
@@ A.1 Binless estimation methods using samples
(and the same for $q_{\mathrm{B}}^{(k)}$), where $\{\tilde{\varphi}^{(k)}_{{\mathrm{A}}n}\}$ are the instantaneous phase samples of agent $\mathrm{A}$}, $\{\varphi^{(k)}_{{\mathrm{A}}n}\}$, arranged in increasing order on a ring ($\tilde{\varphi}^{(k)}_{{\mathrm{A}}n+1}>\tilde{\varphi}^{(k)}_{{\mathrm{A}}n}$, with a $\mathrm{mod}\,1$ constraint) and $1/N$ ensures normalization. In Sec. IV, where a statistically significant estimate of $q^{(k)}(\varphi)$ was needed, we used, rather, the kernel-density estimate of Eq. (11).
@@
There are two technical limitations involving the above binless estimation methods. First is our limited sample size $N=500$, leading to an errorbar of $\pm 0.2$ for entropy values (identified from samples of uniform distribution) and $\pm 0.05$ for KLD values (identified from two sets of samples of the same Gaussian distribution). Second is the limited machine precision [24]; the phases are distinguishable only up to $\Delta_{{\mathrm{A}}n\leftarrow{\mathrm{B}}}^{(k)}>2^{-52}$ using the conventional $64\mathrm{bit}$ representation of real numbers. Each sample with $\Delta_{{\mathrm{A}}n\leftarrow{\mathrm{B}}}^{(k)}<10\times 2^{-52}$ is ignored, and we replace $N\to N-1$.
The loss of samples to machine precision is a prevalent limitation. In the long realization ($k=10^{4}$) of Sec. IV, $407$ and $439$ samples remained in the end for ensembles $\mathrm{A}$} and $\mathrm{B}$}, respectively, which means that our estimates are still statistically meaningful. On the other hand, this implies that the low-entropy distributions (which are those that usually culminate in loss of samples) did not appear much; see Fig. 5, where we see no distribution with S<21S<-21. Upon continuing this simulation, there were no longer any distinguishable phase samples past $k=1.2\cdot 10^{4}$, at which point the dynamics have attained entropies between 34<S<21-34<S<-21. Were lower-entropy distributions to appear, our results of the main text would improve further, as they are particularly useful for having a small discrepancy $\Delta\varphi_{\mathrm{f}}\sim e^{S}/\sqrt{N}$.